OpenStudy (anonymous):

solve show work. -3(7x-6)-x=26-2(11x-4) help pleez

OpenStudy (anonymous):

-21x+18-x = 26-22x+8 -21x-x+22x = 26+8-18 invalid cause x's will cancel out. are you sure you typed right?

OpenStudy (anonymous):

that's wat i got